# Exotic Stable Branches with Efficient TOV Sequences

### v1

@@ Abstract
Modern inference schemes for the neutron star equation of state (EoS) require large numbers of stellar models constructed with different EoS, and these stellar models must capture all the behavior of stable stars. I introduce termination conditions for sequences of stellar models for cold, non-rotating neutron stars that are guaranteed to identify all stable stellar configurations up to arbitrarily large central pressures along with an efficient algorithm to build accurate interpolators for macroscopic properties. I explore the behavior of stars with both high- and low-central pressures. Interestingly, I find that EoS with monotonically increasing sound-speed can produce multiple stable branches (twin stars) and that large phase transitions at high densities can produce stable branches at nearly any mass scale, including sub-solar masses, while still supporting stars with $M>2\,M_{\odot}$. I conclude with some speculation about the astrophysical implications of this behavior.
I Introduction
@@
Neutron stars (NSs) are extremely dense stellar remnants typically observed with masses between 1–2 $M_{\odot}$ with radii $\sim 12\,\mathrm{km}$ [1]. As such, NSs reach characteristic densities [^1] I write expressions in units where $G=c=1$ but provide explicit units when relevant. $\rho\right>=3M/4\pi R^{3}\approx 5\cdot 10^{14}\,\mathrm{g}/\mathrm{cm}^{3}$, roughly twice the density of atomic nuclei (nuclear saturation density: $n_{\mathrm{sat}}\approx 0.16\,\mathrm{fm}^{-3}$ or $\rho_{\mathrm{sat}}=m_{n}n_{\mathrm{sat}}\approx 2.8\times 10^{14}\,\mathrm{g}/\mathrm{cm}^{3}$ where $m_{n}$ is the nucleon rest-mass). Therefore, many-body nuclear interactions play a key role in NS structure. At the same time, NSs can reach compactness $C=M/R\sim 1/4$, which is close to the upper limit expected from a Schwarzschild Black Hole (BH): $C_{\mathrm{Sch}}=1/2$. Clearly, then, relativistic effects should also be significant.
@@
The structure of cold, non-rotating NSs is governed by the Tolman-Oppenheimer-Volkoff (TOV) equations [2, 3] [^2] It is common practice to solve the TOV equations in a slightly different form, sometimes called the (log)enthalpy formulation [4]. See also Ref. [5] for a recent review of numerical techniques. I find that the enthalpy formulation provides a factor of $\sim$ 5-10 speedup compared to the “standard” formulation described above.
@@
The only remaining free parameter is the central pressure ($p_{c}$), and one can construct relationships between, e.g., the radius ($R$) and gravitational mass ($M=m(R)$) by repeatedly solving for a sequence of $p_{c}$. One is often instructed to continue this process until a local maximum in $M$ is observed, which corresponds to a loss of stability [6, 7, 8, 9]. Because only stable stellar configurations can persist in nature, the sequence of stellar models is terminated at that point.
Figure 1: EoS based on chiral effective field theory ($\chi$EFT [10]) below $p=2.2\,\mathrm{MeV}/\mathrm{fm}^{3}$ (top) without and (bottom) with 1st-order phase transitions. I show (left) $M$-$R$, (middle) $M$-$\Lambda$ (tidal deformability), and (right) $M$-$p_{c}$ curves. Stable branches are dark solid lines and unstable branches are light lines of the same color. The points where stability changes are: (circles) even modes ($\mathrm{n}=0,2,...$) and (triangles) odd modes ($\mathrm{n}=1,3,...$) with (filled symbols) where stability is regained and (empty symbols) where stability is lost. For clarity, I only show modes with $\mathrm{n}\leq 2$. I also show the Schwarzschild ($R_{\mathrm{Sch}}=2M$), Buchdahl ($R_{\mathrm{Bch}}=9M/4$), and light ring ($R_{\mathrm{LR}}=3M$) radii along with the approximate 90% credible region for the mass of J0740+6620 [11, 12].
The fact that cold, non-rotating NSs are bounded below a maximum mass ($M_{\mathrm{TOV}}$) is often referred to as the TOV limit, and the stable stellar models that exist with $dM/dp_{c}>0$ and $p_{c}$ less than the central pressure at the TOV limit ($p_{\mathrm{TOV}}$) are called a stable branch of the, e.g., $M$-$R$ curve.
However, in general, the EoS can support multiple stable branches. Famously, EoS with strong 1st-order phase transitions can yield stellar sequences which lose stability when $p_{c}$ reaches the onset of the phase transition and regain stability when $p_{c}$ is slightly above the end of the phase transition [13]. This can lead to multiple stars with the same $M$ but different $R$ (twin stars). Fig. 1 shows a few examples. See, e.g., Refs. [13, 14, 15, 16] for more discussion.
Importantly, the naive termination condition for the TOV sequence fails to capture the physics of phase transitions (i.e., the presence of twin stars). It is clear, then, that one must continue to solve the TOV equations for $p_{c}$ above the first point where $dM/dp_{c}=0$. This begs the question of when, if ever, one can be confident that they have reached the end of the final stable branch. This letter provides a new termination condition based on the stability of high-density stars, which automatically detects the beginning and the end of all stable NS branches.
@@ II TOV termination conditions
through a counterclockwise spiral in the $M$-$R$ plane. [^4] Similar behavior can also appear in Newtonian gravity [17, 18]. The location of the spiral depends on the EoS, but it always contains many local extrema in $M$. The relevant “static” stability criteria [6, 7, 8, 9] can be summarized as follows: a $M$-$R$ curve that bends counterclockwise corresponds to a mode losing stability, while one that bends clockwise corresponds to a mode regaining stability. As discussed extensively in Harrison et al. [7], the fact that $dM/dp_{c}>0$ for some parts of the spiral does not guarantee stability, and instead each subsequent extremum of $M$ within the spiral corresponds to another stellar eigenmode becoming unstable in order of increasing number of radial nodes ($\mathrm{n}$). Harrison et al. [7] even derives analytic expressions for the $M$-$p_{c}$ curve when $c_{s}^{2}$ is a constant at large $p$: $M$ is a damped-sinusoid in $\ln p_{c}$, and there is an infinite sequence of branches where $dM/dp_{c}>0$ above $p_{\mathrm{TOV}}$.
At very large $p_{c}$, I find that once the first harmonic ($\mathrm{n}=1$) becomes unstable, the star can never become completely stable again. The $\mathrm{n}=1$ mode may momentarily restabilize (panel E of Fig. 1), but the fundamental mode ($\mathrm{n}=0$) will not restabilize even if the EoS is causal ($c_{s}^{2}=1$).
In the presence of twin stars, we may see several extrema in $M$ corresponding to $\mathrm{n}=0$ losing and regaining stability. Nevertheless, after $\mathrm{n}=0$ becomes unstable “for good,” the next mode to become unstable is $\mathrm{n}=1$. This corresponds to a local minimum in $M$ with $dR/dp_{c}>0$. At this point, additional extrema never correspond to $\mathrm{n}=0$ regaining stability. Said another way, while $\mathrm{n}=0$ can regain stability at high $p_{c}$ when $\mathrm{n}=1$ is still stable, it cannot once $\mathrm{n}=1$ is unstable. If one detects an unstable $\mathrm{n}=1$, then, this signals the final spiral towards $M_{\circlearrowleft}$ and that all stars with larger $p_{c}$ are unstable.
@@
At low $p_{c}$, one also expects a minimum NS mass associated with $\mathrm{n}=0$ regaining stability after the white dwarf (WD) branch loses stability [7]. This manifests as a local minimum in $M$ with $dR/dp_{c}<0$. Typically, this occurs for $p_{c}$ low enough to be in the NS crust (i.e., the whole star is “crust”), where the physics is relatively well understood and one does not have to worry about exotic behavior.
III Adaptive TOV sequences
@@
After the recursive routine terminates, the algorithm checks whether the termination conditions at high $p_{c}$ ($\mathrm{n}=1$ mode loses stability) and low $p_{c}$ ($\mathrm{n}=0$ mode regains stability) are observed within the existing set of stellar models. If not, it extends the range of $p_{c}$ geometrically as needed and calls the recursive routine on the new segments. This is repeated until the termination conditions are observed (or the algorithm reaches values of $p_{c}$ larger than any $p$ recorded within the tabulated EoS).
@@
This issue does not arise when extending the sequence to higher $p_{c}$ as long as one focuses on the range of $p_{c}$ relevant for NSs and stays well above what is relevant for WDs ($p_{c}\sim 6\times 10^{-7}\,\mathrm{MeV}/\mathrm{fm}^{3}$).
@@
Altogether, my implementation [19] requires $O(3)$ ms/model on a modern laptop (11th Gen Intel i7-1185G7 clocked at 333\,3GHz). Reliable interpolators for $M$-$R$ can be constructed with with $\sim$100-150 models (interpolators for $M$-$\Lambda$ require more points; see discussion below), yielding an runtime of $\lesssim O(500)$ ms/EoS. [^7] It takes longer to read the EoS table from and write the TOV solutions to disk.
@@
Figure 2: EoS with monotonically increasing $c_{s}^{2}$ that produce twin stars: (left) $c_{s}^{2}$ vs. $\varepsilon$ and (right) $M$ vs. $R$. Stable and unstable branches are denoted as in Fig. 1. The ability of such EoS to support twin stars is a combination of (top) the smoothness of $c_{s}^{2}$ and (bottom) the size of the density range where $c_{s}^{2}$ is nearly constant. This behavior is reminiscent of $2^{\mathrm{nd}}$-order phase transitions as there is a (near) discontinuity in $dc_{s}^{2}/d\varepsilon=d^{2}p/d\varepsilon^{2}$ rather than $c_{s}^{2}=dp/d\varepsilon$. Although not shown here, EoS that support twin stars do so even if $c_{s}^{2}\rightarrow 1/3$ at higher densities. I use examples with $M_{\mathrm{TOV}}<2\,M_{\odot}$ to more clearly show the variety of behavior. Fig. 1 shows an example with $M_{\mathrm{TOV}}\sim 2\,M_{\odot}$.
IV Exotic low-density behavior
As briefly discussed before, the EoS can support multiple stable branches at low $p_{c}$ while still matching ab initio theory [10] at $n\lesssim n_{\mathrm{sat}}$. In particular, there can be a phase transition just above $n_{\mathrm{sat}}$ that is consistent with astrophysical observations, which currently allow phase transitions (and multiple stable branches) at $M\lesssim M_{\odot}$ and/or $M\gtrsim 2\,M_{\odot}$ [1, 13]. In fact, although far from being certain, if a large slope parameter at $n_{\mathrm{sat}}$ ($L\gtrsim 100\,\mathrm{MeV}$) is observed experimentally, such a phase transition just above $n_{\mathrm{sat}}$ could actually be preferred by current observations [20, 21].
@@ IV Exotic low-density behavior
V Exotic high-density behavior
Fig. 1 also shows various high-density behavior, including stiff ($c_{s}^{2}\sim 1$) and soft ($c_{s}^{2}\sim 1/3$) EoS both with and without $1^{\mathrm{st}}$-order phase transitions. All stellar sequences eventually end in a counterclockwise spiral. However, the path by which the high-density models enter this spiral can vary. I focus on EoS that approximately satisfy current observations constraints (i.e., follow $\chi$EFT at low densities and have $R(1.4\,M_{\odot})\sim 12.5\,\mathrm{km}$, $M_{\mathrm{TOV}}\gtrsim 2\,M_{\odot}$).
The vanilla behavior is as follows: $\mathrm{n}=0$ becomes unstable, then $M$ decreases and enters a spiral. The spiral is always centered on $M_{\circlearrowleft}<M_{\mathrm{TOV}}$, and the $M$-$R$ curves never cross themselves (i.e., identical twin stars are not physical). It is interesting that EoS with $c_{s}^{2}\sim 1$ for $p\sim p_{\mathrm{TOV}}$ produce $R_{\mathrm{TOV}}$ comparable to (and smaller than!) their light-ring ($R_{\mathrm{LR}}=3M$). This could create a (small) resonating cavity like those invoked to motivate searches for gravitational-wave echoes [22, 23].
The presence of twin stars only moderately complicates this behavior: $\mathrm{n}=0$ can regain stability, and the second stable branch can, but does not have to, reach higher $M$ than the first stable branch. Again, later stable branches with $c_{s}^{2}\sim 1$ can terminate near $R_{\mathrm{LR}}$.

### v2

@@ Abstract
Modern inference schemes for the neutron star equation of state (EoS) require large numbers of stellar models constructed with different EoS, and these stellar models must capture all the behavior of stable stars. I introduce termination conditions for sequences of stellar models for cold, non-rotating neutron stars that can identify all stable stellar configurations up to arbitrarily large central pressures along with an efficient algorithm to build accurate interpolators for macroscopic properties. I explore the behavior of stars with both high- and low-central pressures. Interestingly, I find that EoS with monotonically increasing sound-speed can produce multiple stable branches (twin stars) and that large phase transitions at high densities can produce stable branches at nearly any mass scale, including sub-solar masses, while still supporting stars with $M>2\,M_{\odot}$. I conclude with some speculation about the astrophysical implications of this behavior.
E-mail: ]essick@cita.utoronto.ca
1 Introduction
@@
Neutron stars (NSs) are extremely dense stellar remnants typically observed with masses between 1–2 $M_{\odot}$ with radii $\sim 12\,\mathrm{km}$ (Legred et al., 2021). As such, NSs reach characteristic densities [^1] I write expressions in units where $G=c=1$ but provide explicit units when relevant. $\rho\right>=3M/4\pi R^{3}\approx 5\cdot 10^{14}\,\mathrm{g}/\mathrm{cm}^{3}$, roughly twice the density of atomic nuclei (nuclear saturation density: $n_{\mathrm{sat}}\approx 0.16\,\mathrm{fm}^{-3}$ or $\rho_{\mathrm{sat}}=m_{n}n_{\mathrm{sat}}\approx 2.8\times 10^{14}\,\mathrm{g}/\mathrm{cm}^{3}$ where $m_{n}$ is the nucleon rest-mass). Therefore, many-body nuclear interactions play a key role in NS structure. At the same time, NSs can reach compactness $C=M/R\sim 1/4$, which is close to the upper limit expected from a Schwarzschild Black Hole (BH): $C_{\mathrm{Sch}}=1/2$. Additionally, the speed of sound ($c_{s}$) within the NS core almost certainly exceeds $1/\sqrt{3}$ (Legred et al., 2021) and therefore reaches an appreciable fraction of the speed of light. Clearly, then, relativistic effects should also be significant.
Figure 1: EoS based on chiral effective field theory ($\chi$EFT; Keller et al., 2023) below $p=2.2\,\mathrm{MeV}/\mathrm{fm}^{3}$ and a BPS crust (Baym et al., 1971) (top) without and (bottom) with 1st-order phase transitions. I show (left) $M$-$R$, (middle) $M$-$\Lambda$ (tidal deformability), and (right) $M$-$p_{c}$ curves. Stable branches are dark solid lines and unstable branches are light lines of the same color. The points where stability changes are: (circles) even modes ($\mathrm{n}=0,2,...$) and (triangles) odd modes ($\mathrm{n}=1,3,...$) with (filled symbols) where stability is regained and (empty symbols) where stability is lost. For clarity, I only show modes with $\mathrm{n}\leq 2$. I also show the Schwarzschild ($R_{\mathrm{Sch}}=2M$), Buchdahl ($R_{\mathrm{Bch}}=9M/4$), and light ring ($R_{\mathrm{LR}}=3M$) radii along with the approximate 90% credible region for the mass of J0740+6620 (Cromartie et al., 2019; Fonseca et al., 2021).
@@
The structure of cold, non-rotating NSs is governed by the Tolman-Oppenheimer-Volkoff (TOV) equations (Oppenheimer & Volkoff, 1939; Tolman, 1939) [^2] It is common practice to solve the TOV equations in a slightly different form, sometimes called the (log)enthalpy formulation (Lindblom, 2014). See also Kastaun & Ohme (2024) for a recent review of numerical techniques. I find that the enthalpy formulation provides a factor of $\sim$ 5-10 speedup compared to the “standard” formulation described above.
@@
The only remaining free parameter is the central pressure ($p_{c}$), and one can construct relationships between, e.g., the radius ($R$) and gravitational mass ($M=m(R)$) by repeatedly solving for a sequence of $p_{c}$. One is often instructed to continue this process until a local maximum in $M$ is observed, which corresponds to a loss of stability (Bardeen et al., 1966; Harrison et al., 1965; Sorkin, 1981, 1982). Because only stable stellar configurations can persist in nature, the sequence of stellar models is terminated at that point.
The fact that cold, non-rotating NSs are bounded below a maximum mass ($M_{\mathrm{TOV}}$) is often referred to as the TOV limit, and the stable stellar models that exist with $dM/dp_{c}>0$ and $p_{c}$ less than the central pressure at the TOV limit ($p_{\mathrm{TOV}}$) are called a stable branch of the $M$-$R$ curve.
However, in general, the EoS can support multiple stable branches. Famously, EoS with strong phase transitions can yield stellar sequences which lose stability when $p_{c}$ reaches the onset of the phase transition and regain stability when $p_{c}$ is slightly above the end of the phase transition (Essick et al., 2023). This can lead to multiple stars with the same $M$ but different $R$ (twin stars). Fig. 1 shows a few examples. See, e.g., Essick et al. (2023); Alford et al. (2005, 2013) and Mroczek et al. (2023) for more discussion.
Importantly, the naive termination condition for the TOV sequence fails to capture the physics of phase transitions (i.e., the presence of twin stars). It is clear, then, that one must continue to solve the TOV equations for $p_{c}$ above the first point where $dM/dp_{c}=0$. This begs the question of when, if ever, one can be confident that they have reached the end of the final stable branch. This letter provides such a termination condition based on the stability of high-density stars: i.e., NSs with $p_{c}$ larger than the $p_{c}$ at which the $dM/dp_{c}=0$ and $dR/dp_{c}>0$ (n=1 radial mode loses stability) are almost certainly unstable. This criterion automatically detects the end of all stable NS branches for EoS that support massive stars ($M_{\mathrm{TOV}}\gtrsim 2\,M_{\odot}$). [^4] See Appendix A for counter-examples with EoS that do not support massive stars.
@@ 2 TOV termination conditions
through a counterclockwise spiral in the $M$-$R$ plane. [^5] Similar behavior can also appear in Newtonian gravity (Bonnar, 1956; Ebert, 1955). The location of the spiral depends on the EoS, but it always contains many local extrema in $M$. The relevant “static” stability criteria (Bardeen et al., 1966; Harrison et al., 1965; Sorkin, 1981, 1982) can be summarized as follows: a $M$-$R$ curve that bends counterclockwise corresponds to a mode losing stability, while one that bends clockwise corresponds to a mode regaining stability. As discussed extensively in Harrison et al. (1965), the fact that $dM/dp_{c}>0$ for some parts of the spiral does not guarantee stability, and instead each subsequent extremum of $M$ within the spiral corresponds to another stellar eigenmode becoming unstable in order of increasing number of radial nodes ($\mathrm{n}$). Harrison et al. (1965) even derives analytic expressions for the $M$-$p_{c}$ curve when $c_{s}^{2}$ is a constant at large $p$: $M$ is a damped-sinusoid in $\ln p_{c}$, and there is an infinite sequence of branches where $dM/dp_{c}>0$ above $p_{\mathrm{TOV}}$.
@@
For EoS that support massive stars ($M_{\mathrm{TOV}}\gtrsim 2\,M_{\odot}$) and at very large $p_{c}$, I find that once the first harmonic ($\mathrm{n}=1$) becomes unstable, the star can never become completely stable again. The $\mathrm{n}=1$ mode may momentarily restabilize (in particular, this occurs if a high-$p$ phase transition caused $\mathrm{n}=1$ to lose stability and $c_{s}^{2}=1$ above the transition; panel E of Fig. 1), but the fundamental mode ($\mathrm{n}=0$) will not restabilize even if the EoS is causal ($c_{s}^{2}=1$). Fig. 1 shows several such extensions. The red curve in the top row extends the orange curve with $c_{s}^{2}=1$ at pressures above the $p_{c}$ at which $\mathrm{n}=1$ loses stability ($p_{n=1}$) in the orange curve. The light-blue curve in the bottom row modifies the orange curve from the top row by introducing a $1^{\mathrm{st}}$-order phase transition (which causes $\mathrm{n}=1$ to lose stability) and has $c_{s}^{2}=1$ at higher densities (which causes $\mathrm{n}=1$ to regain stability, but $\mathrm{n}=0$ remains unstable). Appendix A discusses this in more detail.
In the presence of twin stars, one may see several extrema in $M$ corresponding to $\mathrm{n}=0$ losing and regaining stability. Nevertheless, at $p_{c}$ higher than the $p_{c}$ where $\mathrm{n}=0$ becomes unstable “for good,” the next mode to become unstable is $\mathrm{n}=1$. This corresponds to a local minimum in $M$ with $dR/dp_{c}>0$. At this point, additional extrema never correspond to $\mathrm{n}=0$ regaining stability. Said another way, while $\mathrm{n}=0$ can regain stability at high $p_{c}$ when $\mathrm{n}=1$ is still stable, it cannot once $\mathrm{n}=1$ is unstable for massive NSs. If one detects an unstable $\mathrm{n}=1$, then, this signals the final spiral towards $M_{\circlearrowleft}$ and that all stars with larger $p_{c}$ are unstable.
This criterion is motivated by a survey of possible high-density extensions to reasonable EoS (Fig. 1). I do not know of a general proof, and counterexamples exist for low-$C$ stars: e.g., white dwarfs (WD; Baym et al., 1971), strange dwarfs (Alford et al., 2017; Goncalves et al., 2023; Di Clemente et al., 2024), and EoS with $M_{\mathrm{TOV}}\lesssim 1.0\,M_{\odot}$ (Appendix A). The criterion is physically reasonable, though, as restabilization at $p_{c}\geq p_{n=1}$ is only possible if $c_{s}^{2}$ increases dramatically and is associated with an increase in $C$. EoS with large $M_{\mathrm{TOV}}$ already have large $c_{s}^{2}$, and they are nearly as compact as BHs. Therefore, restabilization at $p_{c}\geq p_{n=1}$ does not seem possible for EoSs with large $M_{\mathrm{TOV}}$ because both $c_{s}^{2}$ and $C$ are bounded from above by causality and BH spacetimes (formation of a trapped surface), respectively.
At low $p_{c}$, one also expects a minimum NS mass associated with $\mathrm{n}=0$ regaining stability after the WD branch loses stability (Harrison et al., 1965). This manifests as a local minimum in $M$ with $dR/dp_{c}<0$. Typically, this minimum occurs at $p_{c}$ low enough to be within the NS crust (i.e., the whole star is “crust”), where the physics is relatively well understood and one does not have to worry about exotic behavior.
3 Adaptive TOV sequences
@@
After the recursive routine terminates, the algorithm checks whether the termination conditions at high $p_{c}$ ($\mathrm{n}=1$ mode loses stability) and low $p_{c}$ ($\mathrm{n}=0$ mode regains stability) are observed within the existing set of stellar models. [^8] Importantly, the algorithm is intended to work for NS branches and relies on the user to seed it with an appropriate initial grid of pressures. If not, it extends the range of $p_{c}$ geometrically as needed and calls the recursive routine on the new segments. This is repeated until the termination conditions are observed (or the algorithm reaches values of $p_{c}$ larger than any $p$ recorded within the tabulated EoS).
@@
This issue does not arise when extending the sequence to higher $p_{c}$ as long as one focuses on the range of $p_{c}$ relevant for NSs ($p_{c}\gtrsim\mathrm{MeV}/\mathrm{fm}^{3}$) and stays well above what is relevant for WDs ($p_{c}\sim 6\times 10^{-7}\,\mathrm{MeV}/\mathrm{fm}^{3}$).
@@
Altogether, my implementation (Essick, 2024) requires $O(3)$ ms/model on a modern laptop (333\,3GHz 11th Gen Intel i7-1185G7). Reliable interpolators for $M$-$R$ can be constructed with with $\sim$100-150 models (interpolators for $M$-$\Lambda$ require more points; see discussion below), yielding an runtime of $\lesssim O(500)$ ms/EoS. [^9] It takes longer to read the EoS table from and write the TOV solutions to disk.
@@
Figure 2: EoS with monotonically increasing $c_{s}^{2}$ that produce twin stars: (left) $c_{s}^{2}$ vs. $\varepsilon$, (center) $p$ vs. $\varepsilon$, and (right) $M$ vs. $R$. Stable and unstable branches are denoted as in Fig. 1. The ability of such EoS to support twin stars is a combination of (top) the smoothness of $c_{s}^{2}$ and (bottom) the size of the density range where $c_{s}^{2}$ is nearly constant. This behavior is reminiscent of $2^{\mathrm{nd}}$-order phase transitions as there is a (near) discontinuity in $dc_{s}^{2}/d\varepsilon=d^{2}p/d\varepsilon^{2}$ rather than $c_{s}^{2}=dp/d\varepsilon$. Although not shown here, EoS that support twin stars do so even if $c_{s}^{2}\rightarrow 1/3$ at higher densities. I use examples with $M_{\mathrm{TOV}}<2\,M_{\odot}$ to more clearly show the variety of behavior. Fig. 1 (panel B) shows an example with $M_{\mathrm{TOV}}\sim 2\,M_{\odot}$.
4 Exotic low-density behavior
As briefly discussed before, the EoS can support multiple stable branches at low $p_{c}$ while still matching ab initio theory (Keller et al., 2023) at $n\lesssim n_{\mathrm{sat}}$, and similar behavior has been reported elsewhere (see, e.g., Category IV EoSs in Christian et al., 2018). In particular, there can be a phase transition just above $n_{\mathrm{sat}}$ that is consistent with astrophysical observations, which currently allow phase transitions (and multiple stable branches) at $M\lesssim M_{\odot}$ and/or $M\gtrsim 2\,M_{\odot}$ (Legred et al., 2021; Essick et al., 2023). In fact, although far from being certain, if a large slope parameter at $n_{\mathrm{sat}}$ ($L\gtrsim 100\,\mathrm{MeV}$) is observed experimentally, such a phase transition just above $n_{\mathrm{sat}}$ could actually be preferred by current observations (Essick et al., 2021b, a).
@@ 4 Exotic low-density behavior
5 Exotic high-density behavior
Figure 3: EoS with high-density phase transitions and HiPO branches: (left) $p$ vs. $\varepsilon$, (center) $M$ vs. $R$, and (right) $\Lambda$ vs $M$. Stability is indicated as in Figs. 1 and 2. I show (solid lines) EoS with $c_{s}^{2}=1$ up to arbitrarily high densities and (dashed lines) EoS that match (purple square) approximate pQCD predictions at $\varepsilon\approx 12.5\,\mathrm{GeV}/\mathrm{fm}^{3}$ for $p$ and $c_{s}^{2}$ (Komoltsev et al., 2023). The single EoS that fails to support a HiPO branch does so because $c_{s}^{2}=1/3$ at all densities above its phase transition (i.e., $c_{s}^{2}$ never reaches $\sim 1$). This is also the only EoS which fails to match the pQCD prediction, suggesting that pQCD places an effective lower limit on HiPO masses of $\gtrsim 0.5\,M_{\odot}$.
Fig. 1 also shows various high-density behavior, including stiff ($c_{s}^{2}\sim 1$) and soft ($c_{s}^{2}\sim 1/3$) EoS both with and without phase transitions. All stellar sequences eventually end in a counterclockwise spiral. However, the path by which the high-density models enter this spiral can vary. I focus on EoS that approximately satisfy current observations constraints (i.e., follow $\chi$EFT at low densities and have $R(1.4\,M_{\odot})\sim 12.5\,\mathrm{km}$, $M_{\mathrm{TOV}}\gtrsim 2\,M_{\odot}$). See Appendix A for discussion of more general behavior.
The vanilla behavior for EoS with $M_{\mathrm{TOV}}\gtrsim 2\,M_{\odot}$ is as follows: $\mathrm{n}=0$ becomes unstable, then $M$ decreases and enters a spiral. The spiral is always centered on $M_{\circlearrowleft}<M_{\mathrm{TOV}}$, and the $M$-$R$ curves never cross themselves (i.e., identical twin stars are not physical). It is interesting that EoS with $c_{s}^{2}\sim 1$ for $p\sim p_{\mathrm{TOV}}$ can produce $R_{\mathrm{TOV}}$ comparable to (and smaller than!) their light-ring ($R_{\mathrm{LR}}=3M$). This could create a (small) resonating cavity like those invoked to motivate searches for gravitational-wave echoes (Longo Micchi et al., 2021; Oshita & Afshordi, 2019).
The presence of twin stars only moderately complicates this behavior: $\mathrm{n}=0$ can regain stability, and the second stable branch can, but does not have to, reach higher $M$ than the first stable branch. Again, later stable branches with $c_{s}^{2}\sim 1$ can terminate near $R_{\mathrm{LR}}$. See Alford et al. (2013) and Christian et al. (2018) for categorization schemes for twin-star branches.